# The Fundamental Limits of Least-Privilege Learning

### v1

@@ 5 Conclusions
The promise of least-privilege learning – to learn feature representations that are useful for a given task but avoid leakage of any information that might be misused and cause harm – is extremely appealing. In this paper, we show that in many realistic settings where there exists inherent uncertainty about the task labels, any representation that provides utility for its intended task must always leak information about attributes other than the task and thus does neither fulfil the least-privilege principle nor provide a better utility-privacy trade-off than existing formal privacy-preserving techniques. Furthermore, we show that, due to a task’s fundamental leakage, even representations that fulfil the least-privilege principle might not prevent any potential harms. These issues apply to any setting in which users share representations instead of raw data records and regardless of the learning technique used to learn the feature mapping that produce these representations.
@@ 2. II)
    Maximal leakage satisfies data processing inequalities in a Markov chain YXZY-X-Z:
@@ 3. III)
    $\mathcal{L}(X\rightarrow Z)$ $=\log\left(\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X)}P(z\mid x)\right)$
    $\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathrm{supp}(Y)}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x,y)\right)$
Maximally Revealing Attributes. We demonstrate that a certain family of attributes is maximally revealing, i.e., achieves the supremum $\sup_{S:~S-X-Z}I_{\infty}(S;Z)$.
@@
A family of maximally revealing attributes $\mathcal{S}^{*}(X)$ is a subset of probability distributions over $\mathbb{S}$, defined as follows. Any attribute $S\in\mathcal{S}^{*}(X)$ is such that for any $x\in\mathbb{X}$ there exists a map $\mathcal{U}_{S}(x):\mathbb{X}\rightarrow 2^{\mathbb{S}}$, and $S=s\in\mathcal{U}_{S}(x)$ if and only if $X=x$.
Lemma 2.
The class of maximally revealing attributes leads to the highest inference gain:
$\sup_{S:~S-X-Z}I_{\infty}(S;Z)=\sup_{S\in\mathcal{S}^{*}(X)}I_{\infty}(S;Z)$ (12)
@@
Denoting by $P_{\tilde{X}}\ll P_{X}$ the relation $\mathrm{supp}(P_{\tilde{X}})\subseteq\mathrm{supp}(P_{X})$, observe that any distribution $P_{\tilde{X}}\ll P_{X}$ can be represented as follows:
$P_{\tilde{X}}(x)=\frac{\sum_{s\in\mathcal{U}_{S^{*}}(x)}P_{S^{*}}(s)}{\sum_{x\in\mathbb{X}}\sum_{s\in\mathcal{U}_{S^{*}}(x)}P_{S^{*}}(s)},$ (13)
for some $S^{*}\in\mathcal{S}^{*}(X)$.
Liao et al. (2019, Eq. 106–109) show that for $P_{\tilde{X}}$ and $S^{*}\in\mathcal{S}^{*}(X)$ defined as in Eq. 13, we have:
$I_{\infty}(S^{*};Z)$ $=\sup_{S:~S-\tilde{X}-Z}I_{\infty}(S;Z)$ (14)
Moreover, Liao et al. (2019, Theorem 2) show that the following property holds:
$\sup_{S:~S-X-Z}I_{\infty}(S;Z)$ $=\sup_{\tilde{X}:~P_{\tilde{X}}\ll P_{X}}\sup_{S:~S-\tilde{X}-Z}I_{\infty}(S;Z)$ (15)
Using the fact that for any $\tilde{X}:~P_{\tilde{X}}\ll P_{X}$ there exists $S^{*}\in\mathcal{S}^{*}(X)$ such that Eq. 13 holds, we combine Eq. 14 and Eq. 15 to get the sought result. ∎
Lemma 3.
Suppose that $|\mathbb{X}|>1$. If an attribute is maximally revealing, $S\in\mathcal{S}^{*}(X)$, then the conditional distribution $P_{S\mid X}$ is non-positive: there exist a $s\in\mathbb{S}$ and $x\in\mathbb{X}$ such that $P(s\mid x)=0$.
@@
By contradiction. Suppose that for all $x\in\mathbb{X}$, we have $P(s\mid x)>0.$ But then, there exist $x^{\prime}\neq x^{\prime\prime}\in\mathbb{X}$ such that $s\in\mathcal{U}_{S}(x^{\prime})$ and $s\in\mathcal{U}_{S}(x^{\prime\prime})$, which contradicts the definition of $\mathcal{S}^{*}(X)$. ∎
Omitted Proofs. Next, we provide the proofs of the formal statements in the main body of the paper.
@@
By construction, we have a Markov chain YXZY-X-Z. Observe that $Y\not\in\mathcal{S}^{*}(X)$ by Lemma 3 and strict positivity (Assumption A). Therefore,
$\sup_{S:~S-X-Z,~S\neq Y}I_{\infty}(S;Z)=\sup_{S\in\mathcal{S}^{*}(X)}I_{\infty}(S;Z)=\sup_{S:~S-X-Z}I_{\infty}(S;Z)=\mathcal{L}(X\rightarrow Z)$ (16)
by Lemma 2 and the definition of maximal leakage. Finally, applying the properties in Lemma 1, we have $I_{\alpha}(Y;Z)\leq\mathcal{L}(Y;Z)\leq\mathcal{L}(X\rightarrow Z)=\sup_{S:~S-X-Z,~S\neq Y}I_{\infty}(S;Z)\leq\gamma$. ∎
@@
Theorem 1 only requires that $Y\not\in\mathcal{S}^{*}(X)$. This holds under weaker assumptions than the strict positivity of the posterior (Assumption A). For instance, it is sufficient that there exist any $x,x^{\prime},y$ such that both $P_{X\mid Y}(x\mid y)>0$ and $P_{X\mid Y}(x^{\prime}\mid y)>0$.
@@
$\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathrm{supp}(Y)}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x,y)\right)$ (17)
$=\log\left(\max_{y\in\mathrm{supp}(Y)}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x)\right),$ (18)
where the second equality is by the Markov chain YXZY-X-Z.
@@
$\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathrm{supp}(Y)}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x)\right)$ (19)
$=\log\left(\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X)}P(z\mid x)\right),$ (20)
@@
Finally, to obtain the trade-off, it suffices to observe that $I_{\alpha}(Y;Z)\leq\mathcal{L}(Y\rightarrow Z)\leq\mathcal{L}(X\rightarrow Z)\leq\gamma$, where the second inequality are by properties of maximal leakage in Lemma 1. ∎
One way to interpret this result is that the family of maximally revealing attributes $\mathcal{S}^{*}(X)$ is not sensitive to conditioning on $Y$ so long as $Y\not\in\mathcal{S}^{*}(X)$, which we have by Lemma 3 and strict positivity (Assumption A).
@@
Issa et al. (2019) show that if the map $f_{E}(\cdot)$ is $\varepsilon$-LDP, then $\mathcal{L}(X\rightarrow Z)\leq\varepsilon$, where $Z=f_{E}(X)$. Following the derivation in the proof of Theorem 2, specifically, Eq. 19, we have that $\mathcal{L}(X\rightarrow Z\mid Y)\leq\varepsilon$. ∎
@@ B.1 Details of Setup in Section 4
B.2 Across Learning Techniques
@@
B.3 Across MLaaS settings
@@
B.4 Across Hidden Layers
@@
B.5 Across Model Architecture
To show that the trade-off holds regardless of the exact model architecture, we repeat the experiment outlined in Section 4 using the ResNet-18 model architecture from He et al. (2015) implemented by PyTorch. Training batch size is 32, SGD learning rate is 0.01. The adversary is given access to the feature representations at the last hidden layer of the model.
@@ B.6 Across Datasets
Experiment Setup. In each experiment, we select one attribute as the model’s learning task $Y$ and a second attribute as the sensitive attribute $S$ targeted by the adversary. We repeat each experiment $5$ times to capture randomness of our measurements for both the model and adversary, and show average results across all $5$ repetitions. At the start of the experiment, we split the data into the three sets $D_{T}$, $D_{E}$, and $D_{A}$. We train a TabNet model on the train set $D_{T}$ for the chosen learning task and then estimate the model’s utility on the evaluation set $D_{E}$. We measure the utility of the model-generated representations as the multiplicative gain $\smash{\tilde{I}_{\infty}}(Y;Z)$. After model training and evaluation, we train both the label-only and features adversary on the auxiliary data $D_{A}$. The features adversary is given access to a record’s representation at the last encoding layer of the TabNet encoder (see Arik & Pfister (2021) for details of the model architecture). For a given sensitive attribute $S$, we estimate the adversary’s gain as $\smash{\tilde{I}_{\infty}}(S,Z\mid Y)$.

### v2

@@ 6 Conclusions
The promise of least-privilege learning — to learn feature representations that are useful for a given task but avoid leakage of any information that might be misused and cause harm — is extremely appealing. In this paper, we show that in certain realistic settings where there exists inherent uncertainty about the task labels, any representation that provides utility for its intended task must always leak information about attributes other than the task and thus does neither fulfil the least-privilege principle nor provide a better utility-privacy trade-off than existing formal privacy-preserving techniques. Furthermore, we show that, due to a task’s fundamental leakage, even representations that fulfil the least-privilege principle might not prevent all the potential harms. These issues apply to any setting in which users share representations instead of raw data records and regardless of the learning technique used to learn the feature mapping that produce these representations.
Impact Statement
This paper validates claims from prior works that the concept of least-privilege learning might provide a promising avenue to mitigate harms of potential misuse of machine learning. Prior works repeatedly suggest that least-privilege learning might achieve a better trade-off between revealing information about an intended task and unintended information leakage than existing techniques, such as differential privacy. In this paper, we formally prove that least-privilege learning might not fulfil this promise: We show that there is an inherent trade-off between a representation’s utility and its unintended information leakage that cannot be overcome by any learning technique. The results of this paper will help practitioners to evaluate the potential benefits and shortcomings of a newly proposed technique. Our formalisation of a previously only informally stated property will enable researchers to formally compare the trade-offs of alternative methods to prevent data misuse in machine learning.
Acknowledgements
This work is partially funded by the Swiss National Science Foundation under grant 200021-188824. NP acknowledges support from the Canada CIFAR AI Chair. We would like to thank Christian Knabenhans for providing a basis for the implementation of the empirical evaluation of this work.
@@ 2. II)
    Maximal leakage satisfies data processing inequalities in a Markov chain $Y-X-Z$:
@@ 3. III)
    $\mathcal{L}(X\rightarrow Z)$ $=\log\left(\sum_{z\in\mathbb{Z}}\max_{x\in\mathbb{X}}P(z\mid x)\right)$
    $\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathbb{Y}}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x,y)\right)$
Maximally Revealing Attributes. We demonstrate that there exists an attribute that is maximally revealing, i.e., achieves the supremum $\sup_{S:~{}S-X-Z}I_{\infty}(S;Z)$, and show how to construct this attribute.
@@
For a given $P_{X}$, a maximally revealing attribute $S^{*}$ (see “shattering distribution”, Issa et al., 2019) is defined over $\mathbb{S}~{}\triangleq~{}\mathbb{X}\times\mathbb{Z}$ via the following conditional probability distribution:
$P_{S^{*}\mid X}((x^{\prime},k)\mid x)~{}\triangleq~{}\begin{cases}\frac{p_{\min}}{P_{X}(x)},&x^{\prime}=x\text{ and }1\leq k\leq\lfloor r(x)\rfloor\\ 1-\frac{(\lceil r(x)\rceil-1)\cdot p_{\min}}{P_{X}(x)},&x^{\prime}=x\text{ and}k=\lceil r(x)\rceil\\ 0,&x^{\prime}\neq x,\end{cases}$ (12)
for any $x,x^{\prime}\in\mathbb{X}$, and $k\in[1,\ldots,\lceil r(x)\rceil]$, where $p_{\min}~{}\triangleq~{}\min_{x\in\mathbb{X}}P_{X}(x)$ and $r(x)~{}\triangleq~{}\nicefrac{{P_{X}(x)}}{{p_{\min}}}$.
Proposition 3 (Issa et al. (2019)).
For any $Z=f_{E}(X)$ and $P_{X}$, the maximally revealing attribute $S^{*}$ leads to the highest inference gain:
$\sup_{S:~{}S-X-Z}I_{\infty}(S;Z)=I_{\infty}(S^{*};Z)$ (13)
@@
It is sufficient to show that for any $P_{Z\mid X}$ and $P_{X}$, we have $\mathcal{L}(X\rightarrow Z)=I_{\infty}(S^{*};Z)$. By Lemma 1 and the definition of $I_{\infty}$, this is equivalent to showing:
$\log\left(\sum_{z\in\mathbb{Z}}\max_{x\in\mathbb{X}}P_{Z\mid X}(z\mid x)\right)=\log\left(\frac{\sum_{z\in\mathbb{Z}}\max_{s\in\mathbb{S}}P_{S^{*},Z}(s,z)}{\max_{s\in\mathbb{S}}P_{S^{*}}(s)}\right).$ (14)
We can see that $\max_{s\in\mathbb{S}}P_{S^{*}\mid Z}(s,z)$ from the RHS has the following form:
$\max_{(x^{\prime},k)\in\mathbb{S}}P_{S^{*},Z}((x^{\prime},k),z)$ $=\max_{(x^{\prime},k)\in\mathbb{S}}P_{S^{*}\mid X}((x^{\prime},k)\mid x^{\prime})\cdot P_{X}(x^{\prime})\cdot P_{Z\mid X}(z\mid x^{\prime})$ (15)
$=\max_{(x^{\prime},k)\in\mathbb{S}\text{ s.t. }k=1}P_{S^{*}\mid X}((x^{\prime},1)\mid x^{\prime})\cdot P_{X}(x^{\prime})\cdot P_{Z\mid X}(z\mid x^{\prime})$ (16)
$=\max_{x\in\mathbb{X}}\frac{p_{\min}}{P_{X}(x)}\cdot P_{X}(x)\cdot P_{Z\mid X}(z\mid x)=p_{\min}\cdot\max_{x\in\mathbb{X}}P_{Z\mid X}(z\mid x),$ (17)
where the first equality is by Markov chain $S^{*}-X-Z$, and the second and third are by Definition 4. Combining the resulting expression and the fact that by Definition 4 we have
$\max_{s\in\mathbb{S}}P_{S^{*}}(s)=\max_{s\in\mathbb{S}}\sum_{x\in\mathbb{X}}P_{S^{*}\mid X}(s)\cdot P_{X}(x)=p_{\min},$ (18)
we get the sought result. ∎
We also show the following technical property of a maximally revealing attribute.
Proposition 4.
Suppose that $|\mathbb{X}|>1$. The conditional distribution $P_{S^{*}\mid X}$ is non-positive: there exist a $s\in\mathbb{S}$ and $x\in\mathbb{X}$ such that $P_{S^{*}\mid X}(s\mid x)=0$.
@@
By Definition 4, for any $x\in\mathbb{X}$ it is sufficient to take any $(x^{\prime},k)\in\mathbb{S}$ where $x^{\prime}\neq x$.
A.1 Omitted Proofs
Next, we provide the proofs of the formal statements in the main body of the paper.
@@ A.1 Omitted Proofs
By construction, we have a Markov chain $Y-X-Z$. Observe that $S^{*}\neq Y$ by Proposition 4 and strict positivity (Assumption A). Therefore,
$\sup_{S:~{}S-X-Z,~{}S\neq Y}I_{\infty}(S;Z)=I_{\infty}(S^{*};Z)=\sup_{S:~{}S-X-Z}I_{\infty}(S;Z)=\mathcal{L}(X\rightarrow Z)$ (19)
by Proposition 3 and the definition of maximal leakage. Finally, applying the properties in Lemma 1, we have $I_{\alpha}(Y;Z)\leq\mathcal{L}(Y\rightarrow Z)\leq\mathcal{L}(X\rightarrow Z)=\sup_{S:~{}S-X-Z,~{}S\neq Y}I_{\infty}(S;Z)\leq\gamma$. ∎
@@
Theorem 1 only requires that $Y\neq S^{*}$. This holds under weaker assumptions than the strict positivity of the posterior (Assumption A). For instance, it is sufficient that there exist any $x,x^{\prime},y$ such that both $P_{X\mid Y}(x\mid y)>0$ and $P_{X\mid Y}(x^{\prime}\mid y)>0$.
Proof of Proposition 1.
For $\gamma=0$, by Lemma 1 we have that $\mathcal{L}(X\rightarrow Z\mid Y)=I(X;Z\mid Y)=0$. As conditional mutual information is consistent with conditional independence, this holds if and only if we have Markov chain $X-Y-Z$. ∎
Proof of Corollary 1.
If $Y=g(X),$ where $g(\cdot)$ is a deterministic function, we can choose $Z=f(Y)=f(g(X)).$ Evidently, this satisfies the Markov relationship $f(g(X))-g(X)-X,$ i.e., $Z-Y-X.$ (Note that $f(\cdot)$ may even be a randomized function for this to hold.) But this Markov relation implies $\mathcal{L}(X\rightarrow Z\mid Y)=0.$
Corollary 2 follows as a consequence of the conditional independence in Proposition 1 and the fact that we have $Y-X-Z$ by construction.
@@
$\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathbb{Y})}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x,y)\right)$ (20)
$=\log\left(\max_{y\in\mathbb{Y}}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x)\right),$ (21)
where the second equality is by the Markov chain $Y-X-Z$.
@@
$\mathcal{L}(X\rightarrow Z\mid Y)$ $=\log\left(\max_{y\in\mathbb{Y}}\sum_{z\in\mathbb{Z}}\max_{x\in\mathrm{supp}(X\mid Y=y)}P(z\mid x)\right)$ (22)
$=\log\left(\sum_{z\in\mathbb{Z}}\max_{x\in\mathbb{X}}P(z\mid x)\right),$ (23)
@@
Finally, to obtain the trade-off, it suffices to observe that $I_{\alpha}(Y;Z)\leq\mathcal{L}(Y\rightarrow Z)\leq\mathcal{L}(X\rightarrow Z)\leq\gamma$, where the the first and the second inequalities are by properties of maximal leakage in Lemma 1. ∎
One way to interpret this result is that the maximally revealing attribute $S^{*}$ is not sensitive to conditioning on $Y$ so long as we have strict positivity of the posterior (Assumption A).
@@
Issa et al. (2019) show that if the map $f_{E}(\cdot)$ is $\varepsilon$-LDP, then $\mathcal{L}(X\rightarrow Z)\leq\varepsilon$, where $Z=f_{E}(X)$. From the form in Eq. 22, we have that $\mathcal{L}(X\rightarrow Z\mid Y)\leq\mathcal{L}(X\rightarrow Z)\leq\varepsilon$. ∎
@@ B.1 Details of Setup in Section 4
B.2 Adversarial Inference Gain
[…]
Fig. 5 shows the adversary’s inference gain $I_{\infty}(S;Z\mid Y)$ which measures the adversary’s classification accuracy for sensitive attribute $S$ (each column) normalised by a task’s (each row) fundamental leakage.
B.3 Across Learning Techniques
@@
B.4 Across MLaaS settings
@@
B.5 Across Hidden Layers
@@
B.6 Across Model Architecture
To show that the trade-off holds regardless of the exact model architecture, we repeat the experiment outlined in Section 4 using the ResNet-18 model architecture from (He et al., 2015) implemented by PyTorch. Training batch size is 32, SGD learning rate is 0.01. The adversary is given access to the feature representations at the last hidden layer of the model.
@@ B.7 Across Datasets
Experiment Setup. In each experiment, we select one attribute as the model’s learning task $Y$ and a second attribute as the sensitive attribute $S$ targeted by the adversary. We repeat each experiment $5$ times to capture randomness of our measurements for both the model and adversary, and show average results across all $5$ repetitions. At the start of the experiment, we split the data into the three sets $D_{T}$, $D_{E}$, and $D_{A}$. We train a TabNet model on the train set $D_{T}$ for the chosen learning task and then estimate the model’s utility on the evaluation set $D_{E}$. We measure the utility of the model-generated representations as the multiplicative gain $\smash{\tilde{I}_{\infty}}(Y;Z)$. After model training and evaluation, we train both the label-only and features adversary on the auxiliary data $D_{A}$. The features adversary is given access to a record’s representation at the last encoding layer of the TabNet encoder (see (Arik & Pfister, 2021) for details of the model architecture). For a given sensitive attribute $S$, we estimate the adversary’s gain as $\smash{\tilde{I}_{\infty}}(S,Z\mid Y)$.